# V2P: From Background Suppression to Center Peaking for Robust GUI Grounding Task

### v1

@@ Abstract
Precise localization of GUI elements is crucial for the development of GUI agents. Traditional methods rely on bounding box or center-point regression, neglecting spatial interaction uncertainty and visual-semantic hierarchies. Recent methods incorporate attention mechanisms but still face two key issues: (1) ignoring processing background regions causes attention drift from the desired area, and (2) uniform labeling fails to distinguish between center and edges of the target UI element, leading to click imprecision. Inspired by how humans visually process and interact with GUI elements, we propose the Valley-to-Peak (V2P) method to address these issues. To mitigate background distractions, V2P introduces a suppression attention mechanism that minimizes the model’s focus on irrelevant regions to highlight the intended region. For the issue of center-edge distinction, V2P applies a Fitts’ Law-inspired approach by modeling GUI interactions as 2D Gaussian heatmaps where the weight gradually decreases from the center towards the edges. The weight distribution follows a Gaussian function, with the variance determined by the target’s size. Consequently, V2P effectively isolates the target area and teaches the model to concentrate on the most essential point of the UI element. The model trained by V2P achieves the performance with 92.3% and 50.5% on two benchmarks ScreenSpot-v2 and ScreenSpot-Pro. Ablations further confirm each component’s contribution, highlighting V2P’s generalizability for precise GUI grounding tasks.
Figure 1: Performance of different baselines and our V2P-7B model on the dataset of ScreenSpot-v2 (left) and ScreenSpot-Pro (right). By training Qwen2.5-VL-7B using our Valley-to-Peak training strategy, our model achieves the best performance among all competitors.
@@ 1 Introduction
Recent advances in large language models (LLMs) and vision-language models (VLMs) have enabled agents to interpret natural language instructions and interact with graphical user interfaces (GUIs) across desktop, mobile, and web platforms. Central to this capability is GUI grounding—the alignment of language commands with semantically relevant UI elements and their spatial locations Cheng et al. (2024). This task bridges user intent and interface actions, supporting the development of intelligent, general-purpose agents for real-world human-computer interaction.
Early GUI grounding approaches framed the task as regression, outputting a bounding box or $(x,y)$ coordinate for a natural-language query Zhang et al. (2025); Qin et al. (2025). However, this “coordinate generation” method suffers weak spatial–semantic alignment Wu et al. (2025), treating coordinates like ordinary words without inherent spatial meaning. Moreover, point-wise regression contradicts the multi-point validity inherent in real interactions. Recent work addresses these issues by leveraging the model’s attention maps Wu et al. (2025). Instead of predicting coordinates, it extracts cross-modal attention weights linking instruction tokens to image patches, selecting the most attended patch as the click position. This approach offers dense spatial supervision and naturally tolerates multiple valid click regions, aligning better with human behavior.
@@
However, empirical evidence reveals two persistent problems.
@@
    Background distraction: the current methods’ loss functions only require the model to increase attention on the target patches but do not constrain it to maintain low attention on the background patches. As a result, the model may assign high scores to background regions as well, leading to a "divergent" attention distribution. Due to the final softmax normalization, high-scoring background areas can absorb part of the normalized probability mass, which weakens or even shifts the attention peak on the target regions.
@@
    Centre-edge confusion: existing labels treat every pixel inside the bounding box equally, the model trained on such data is unable to differentiate between the center and edge regions of a UI element. As a result, it may allocate attention uniformly across both center and edge regions. This often leads to a prediction error where the click location deviates from the target center (i.e., an inaccurate click). Furthermore, when the labels of two UI elements partially overlap (e.g., a button overlapping with an input field), the uniform labeling scheme may cause the edge regions’ label values to accumulate excessively. This, in turn, can lead the model to overly focus on incorrect (and less relevant) areas while neglecting the truly important target center region.
@@
This naturally raises the question: How can we guide the model’s visual attention to concentrate more precisely on the target UI element? Motivated by how humans first isolate a target (valley suppression) and then focus on its most clickable point (peak emphasis), we present Valley-to-Peak (V2P). V2P tackles the two issues synergistically:
Suppression Attention. To minimize distractions from irrelevant regions (such as background patches, icons, or text), we introduce an inverse attention regularization Li et al. (2018). This regularization penalizes high attention weights outside the target, creating low-attention "valleys" around non-target areas. By isolating the target while filtering out false positives, the network improves its focus on the true actionable UI elements in the interface.
Fitts-Gaussian Peak Modeling. Inspired by Fitts’ Law MacKenzie (1992); Fitts (1954), which formalizes human target acquisition performance as a function of target size and distance, we model the spatial distribution of interaction likelihood using a 2D Gaussian centered on the target element, with variance scaled proportionally to its spatial extent. This creates a probabilistic heatmap that is highest at the element’s core and decays smoothly towards the edges, naturally encoding the tolerance of human interaction. The resulting heatmap captures the spatial variability of interactions and steers the model toward a distinct attention peak at the most probable interaction location, thereby enhancing grounding precision.
Through joint optimization of these components, V2P reshapes the attention landscape, suppressing distractions to form valleys in irrelevant areas while emphasizing a sharp peak at the target’s actionable center. This balanced design significantly enhances grounding precision by aligning the model’s behavior with human interaction patterns.
We evaluate V2P on two public benchmarks: ScreenSpot-v2 Wu et al. (2024b) and the more challenging ScreenSpot-Pro Li et al. (2025). Our method achieves 92.3% on ScreenSpot-v2 and 50.5% on ScreenSpot-Pro, representing relative improvements of 3.5% and 23.7% respectively. Comprehensive ablation studies demonstrate that both Suppression Attention and Fitts-Gaussian Labeling consistently contribute to performance gains, highlighting the broad applicability of V2P to high-precision GUI grounding tasks.
@@ 2.1 GUI-Agents
GUI agents have progressed from rudimentary random- or rule-based test tools to multimodal, LLM-driven systems that can follow natural-language instructions. Early efforts such as Monkey testing Wetzlmaier et al. (2016) and planning or script record-and-replay frameworks Memon et al. (2001); Steven et al. (2000) provided basic coverage but required hand-crafted rules or scripts. Machine-learning techniques later enabled more adaptive behaviour: Humanoid Li et al. (2020) and Deep GUI YazdaniBanafsheDaragh & Malek (2022) learned user-like action policies from screenshots, while widget detectors White et al. (2019) improved element recognition. Natural-language interfaces soon followed, e.g. FLIN Mazumder & Riva (2021) and RUSS Xu et al. (2021), and reinforcement learning environments like WoB Shi et al. (2017) and WebShop Yao et al. (2023) pushed web-scale interaction. The recent arrival of LLMs has unified perception, reasoning and control: WebAgent Gur et al. (2024) and WebGUM Furuta et al. (2024) achieve open-world browsing, AutoDroid Wen et al. (2024) and AppAgent Zhang et al. (2023) automate smartphones, and desktop agents such as UFO Zhang et al. (2024) demonstrate GPT-4-level capabilities; industrial systems (e.g. Claude 3.5 Sonnet and Operator) further attest to the practical traction of GUI agents.
2.2 GUI-Actor
GUI-Actor Wu et al. (2025) is the first VLM–based GUI agent that abandons explicit coordinate generation and instead performs coordinate-free grounding. A dedicated <ACTOR> token is appended to the vision–language encoder; its cross–modal attention to visual patch tokens is interpreted as a heat map from which the click location is taken. To cope with the fact that many pixels inside a UI element are equally valid, GUI-Actor adopts multi-patch supervision: every patch whose centre falls inside the ground-truth bounding box receives an identical positive label, whereas all remaining patches are treated as negatives. This design alleviates the granularity mismatch between patch-level features and pixel-level actions and yields a model that generalises better than previous coordinate-regression baselines across web, mobile and desktop benchmarks. GUI-Actor also introduces a lightweight grounding verifier that reranks the top-k attention peaks, further improving robustness.
Despite these advances, two issues remain. (1) Because the loss encourages high scores only on foreground patches and ignores the background, attention can drift toward visually salient but irrelevant areas. (2) Uniformly labelling all foreground patches fails to distinguish the centre of the element from its edges, which leads to imprecise clicks and error amplification when neighbouring elements overlap. Moreover, the additional verifier introduces a two-stage pipeline and extra inference cost, making the overall design less elegant.
Our Valley-to-Peak (V2P) framework inherits GUI-Actor’s coordinate-free backbone but resolves these issues with a single-stage solution: suppression attention explicitly down-weights backgrounds, and Fitts–Gaussian peak supervision highlights the element centre—eliminating the need for an external verifier while further boosting grounding accuracy.
@@ 3 Method
How can we guide a model’s attention to focus precisely on a target UI element while suppressing distractions? Inspired by human interaction patterns MacKenzie (1992); Fitts (1954), we present Valley-to-Peak (V2P), a framework that reshapes attention landscapes to better reflect real-world behavior through two key components:
  • Suppression Attention Valley Constraint: To reduce distractions, we introduce inverse attention regularization that penalizes high attention weights outside the target, forming low-attention “valleys” around irrelevant regions.
  • Fitts-Gaussian Peak Modeling: Uses a 2D Gaussian centered on the target to model interaction likelihood, creating a peak at the target’s core and guiding attention for precise grounding.
By jointly optimizing suppression and emphasis, V2P refines spatial focus on the actionable center while suppressing distractions. Also, unlike prior methods that treat UI elements rigidly, V2P employs size-adaptive Gaussian supervision to produce continuous, spatially graded attention maps that align with human patterns. The proposed training method significantly improves both the robustness and precision of UI localization, ensuring reliable performance under various challenges.
We describe the overall architecture in Section 3.1, then introduce suppression attention in Section 3.2 for reducing distractions. Finally, Section 3.3 details our Gaussian soft labeling scheme, which better aligns the model’s attention with natural human interaction patterns.
@@ 3.1 Model Architecture Overview
We build upon GUI-Actor Wu et al. (2025), a coordinate-free visual grounding framework that localizes GUI actions through attention rather than coordinate regression. Given a screenshot $I$ and an instruction $q$, the model introduces a special token <ACTOR> in the output sequence as a contextual anchor. The final-layer hidden state of <ACTOR>, denoted $h_{\texttt{<ACTOR>}}$, is used to compute action attention over image patch features $\{v_{1},\dots,v_{M}\}$ extracted by the vision encoder.
@@
$\tilde{v}_{1},\dots,\tilde{v}_{M}=\text{SelfAttn}(v_{1},\dots,v_{M}),$ (1)
yielding contextualized representations. These are projected into a shared embedding space with $h_{\texttt{<ACTOR>}}$ via separate MLPs:
$z$ $=\text{MLP}_{T}(h_{\texttt{<ACTOR>}}),$ (2)
@@
$\alpha_{i}=\frac{z^{\top}z_{i}}{\sqrt{d}},\quad a_{i}=\frac{\exp(\alpha_{i})}{\sum_{j=1}^{M}\exp(\alpha_{j})},$ (4)
@@ 3.2 Suppression Attention Constraint for Distraction Mitigation
$\mathcal{G}=\left\{i\in\{1,\dots,M\}\mid R_{i}\cap b=\emptyset\right\}.$ (5)
We define the suppression attention loss as the total attention mass over these irrelevant regions:
$\mathcal{L}_{\text{Sup\_Attn}}=\sum_{i\in\mathcal{G}}a_{i}.$ (6)
@@ 3.3 Fitts-Gaussian Peak Modeling for Center-Focused Grounding
To better align the model’s attention with human interaction behavior, we introduce Fitts-Gaussian Peak Modeling, a supervision strategy that explicitly encodes the expected attention peak over a target UI element. Our design is grounded in Fitts’ Law MacKenzie (1992); Fitts (1954), which models human pointing behavior: users naturally aim at the center of a target, and larger targets allow for greater spatial tolerance during interaction.
We model the ideal attention distribution as a 2D Gaussian density centered at the centroid of the ground-truth bounding box $b=[x_{1},y_{1},x_{2},y_{2}]$:
$\mu=(c_{x},c_{y})=\left(\frac{x_{1}+x_{2}}{2},\frac{y_{1}+y_{2}}{2}\right).$ (7)
@@
$\sigma_{x}=\frac{w}{\sigma_{\text{factor}}},\quad\sigma_{y}=\frac{h}{\sigma_{\text{factor}}},$ (8)
@@
$y_{i}=\left[\Phi(x^{i}_{\max};c_{x},\sigma_{x})-\Phi(x^{i}_{\min};c_{x},\sigma_{x})\right]\cdot\left[\Phi(y^{i}_{\max};c_{y},\sigma_{y})-\Phi(y^{i}_{\min};c_{y},\sigma_{y})\right],$ (10)
@@
To supervise the model’s predicted attention distribution $\{a_{i}\}$, we adopt the action attention loss from GUI-Actor Wu et al. (2025), using the Kullback-Leibler (KL) divergence to measure the discrepancy between the target $p$ and prediction $a$:
$\mathcal{L}_{\text{Action\_Attn}}=\sum_{i=1}^{M}p_{i}\log\frac{p_{i}}{a_{i}},\quad p_{i}=\frac{y_{i}}{\sum_{j=1}^{M}y_{j}+\epsilon},\quad i=1,\ldots,M,$ (11)

### v3

@@ Abstract
Precise localization of GUI elements is crucial for the development of GUI agents. Traditional methods rely on bounding box or center-point regression, neglecting spatial interaction uncertainty and visual-semantic hierarchies. Recent methods incorporate attention mechanisms but still face two key issues: (1) ignoring processing background regions causes attention drift from the desired area, and (2) uniform modeling the target UI element fails to distinguish between its center and edges, leading to click imprecision. Inspired by how humans visually process and interact with GUI elements, we propose the Valley-to-Peak (V2P) method to address these issues. To mitigate background distractions, V2P introduces a suppression attention mechanism that minimizes the model’s focus on irrelevant regions to highlight the intended region. For the issue of center-edge distinction, V2P applies a Fitts’ Law-inspired approach by modeling GUI interactions as 2D Gaussian heatmaps where the weight gradually decreases from the center towards the edges. The weight distribution follows a Gaussian function, with the variance determined by the target’s size. Consequently, V2P effectively isolates the target area and teaches the model to concentrate on the most essential point of the UI element. The model trained by V2P achieves the performance with 92.4% and 52.5% on two benchmarks ScreenSpot-v2 and ScreenSpot-Pro. Ablations further confirm each component’s contribution, underscoring V2P’s generalizability in precise GUI grounding tasks and its potential for real-world deployment in future GUI agents.
[…]
V2P: Visual Attention Calibration for GUI Grounding via Background Suppression and Center Peaking
Jikai Chen1,2*, Long Chen2*, Dong Wang2*,
[…]
1Zhejiang University  2Inclusion AI, Ant Group
[Uncaptioned image] Model   Code
(a) ScreenSpot-v2
(b) ScreenSpot-Pro
Performance comparison of different baselines and our V2P-7B model on ScreenSpot-v2 (left) and ScreenSpot-Pro (right). By training Qwen2.5-VL-7B using our Valley-to-Peak training strategy, our model achieves the best performance among all competitors.
footnotetext: *Equal contributions. $\ddagger$}Corresponding Authors.
@@ 1 Introduction
Recent advances in large language models (LLMs) and vision-language models (VLMs) have enabled agents to interpret natural language instructions and interact with graphical user interfaces (GUIs) across desktop, mobile, and web platforms. Central to this capability is GUI grounding, which aligns language commands with semantically relevant UI elements and their spatial locations (Cheng et al., 2024). This task bridges user intent and interface actions, supporting the development of intelligent, general-purpose agents for real-world human-computer interaction.
Early approaches framed GUI grounding as coordinate generation task, outputting a bounding box or $(x,y)$ coordinate for a natural-language query (Zhang et al., 2025; Qin et al., 2025). However, this “coordinate generation” method suffers weak spatial–semantic alignment (Wu et al., 2025), treating coordinates like ordinary words without inherent spatial meaning. Moreover, point-wise regression contradicts the multi-point validity inherent in real interactions. Recent work addresses these issues by leveraging the model’s attention maps (Wu et al., 2025). Instead of predicting coordinates, it extracts cross-modal attention weights linking instruction tokens to image patches, selecting the most attended patch as the click position. This approach offers dense spatial supervision and naturally tolerates multiple valid click regions, aligning better with human behavior.
@@
However, after manually scrutinizing the attention heatmap of these methods mentioned above, we found two main issues, as shown in Fig. 1:
@@
    Background Distraction: Current loss functions only reward attention on target patches but fail to penalize it on the background. This leads to a "divergent" attention distribution where background regions also receive high scores. Consequently, softmax normalization allows these regions to absorb probability mass, weakening or even shifting the intended attention peak.
@@
    Centre-edge Confusion: Because labels treat all pixels within a bounding box equally, the model cannot differentiate an element’s center from its edges, resulting in uniform attention and inaccurate clicks that miss the center. Furthermore, for small elements, this often leads the attention to drift towards the edges, making the model more prone to mislocalization, especially when elements overlap.
@@
This raises a key question: How can we guide the model’s attention to focus more precisely on the target UI element? Motivated by human behavior—first isolating the target (valley suppression) then focusing on the action point (peak emphasis)—we propose Valley-to-Peak (V2P). V2P suppresses distractions by creating low-attention "valleys" in irrelevant areas while sharpening a "peak" at the actionable center.
Suppression Attention: We apply inverse attention regularization (Li et al., 2018) to penalize high attention outside the target, isolating true UI elements and reducing attention to non-target regions.
Fitts-Gaussian Peak Modeling: Inspired by Fitts’ Law (MacKenzie, 1992; Fitts, 1954), we use a 2D Gaussian centered on the target, scaled to its size, to model human’s click likelihood, which yields a heatmap that peaks at the center and decays towards the edges, better matching real user interactions.
Together, these modules reshape the attention map, enhancing grounding precision by aligning the model’s focus with human patterns.
On ScreenSpot-v2 (Wu et al., 2024) and the challenging ScreenSpot-Pro (Li et al., 2025), V2P achieves 92.4% and 52.5% element accuracy, significantly outperforming previous methods (see Fig. 1). Ablation studies confirm that both components consistently contribute to performance gains, demonstrating V2P’s broad applicability to high-precision GUI grounding.
Our contribution can be summarized as follows:
  1. 1.
    We systematically analyze existing attention-based methods for visual grounding in GUI agents and, through statistical evaluation, identify two main issues——Background Distraction and Center-Edge Confusion. In addition, we provide a detailed analysis of the underlying causes of these issues and provide insights for further improvements.
  2. 2.
    We introduce Attention Suppression Mechanism (SA) to mitigate Background Distraction and employ Fitts-Gaussian Peak Modeling (FGPM) to effectively alleviate Center-Edge Confusion. Building on these methods, we propose the Valley-to-Peak (V2P) framework, an agentic learning paradigm for GUI grounding that significantly enhances the localization precision and accuracy of Vision-Language Models on GUI elements.
  3. 3.
    Extensive experiments demonstrate that V2P achieves advanced performance on multiple public benchmarks, reaching 92.4% on ScreenSpot-v2 and 52.50% on the challenging ScreenSpot-Pro, with relative improvements of 3.6% and 25.7%. Furthermore, we confirm that V2P demonstrates significant practical value for real-world deployment and seamless integration into GUI agents.
@@ 2.1 GUI-Agents
GUI agents have progressed from rudimentary random- or rule-based test tools to multimodal, LLM-driven systems that can follow natural-language instructions. Early efforts such as Monkey testing (Wetzlmaier et al., 2016) and planning or script record-and-replay frameworks (Memon et al., 2001; Steven et al., 2000) provided basic coverage but required hand-crafted rules or scripts. Machine-learning techniques later enabled more adaptive behaviour: Humanoid (Li et al., 2020) and Deep GUI (YazdaniBanafsheDaragh and Malek, 2022) learned user-like action policies from screenshots, while widget detectors (White et al., 2019) improved element recognition. Natural-language interfaces soon followed, e.g. FLIN (Mazumder and Riva, 2021) and RUSS (Xu et al., 2021), and reinforcement learning environments like WoB (Shi et al., 2017) and WebShop (Yao et al., 2023) pushed web-scale interaction. The recent arrival of LLMs has unified perception, reasoning and control: WebAgent (Gur et al., 2024) and WebGUM (Furuta et al., 2024) achieve open-world browsing, AutoDroid (Wen et al., 2024) and AppAgent (Zhang et al., 2023) automate smartphones, and desktop agents such as UFO (Zhang et al., 2024) demonstrate GPT-4-level capabilities; industrial systems (e.g. Claude 3.5 Sonnet and Operator) further attest to the practical traction of GUI agents.
2.2 GUI Grounding
Prevalent approaches in GUI grounding typically frame the problem as a coordinate generation task (Zhang et al., 2025). Models such as UI-TARS (Qin et al., 2025) and CogAgent (Hong et al., 2024) utilize massive supervised fine-tuning to train VLMs to autoregressively generate textual numerical coordinates to ground the target element. However, treating spatial coordinates as ordinary language tokens can limit fine-grained visual alignment. Consequently, recent methods have largely shifted to leveraging the cross-modal attention maps of Vision-Language Models (VLMs) (Wu et al., 2025). In this paradigm, the model’s prediction is derived from the image patch with the highest attention score in response to a language command. While more robust, this approach often suffers from imprecise attention, with focus leaking into irrelevant background regions or spreading too uniformly across the target element. Our work directly addresses this by refining the quality of the attention map itself.
Our approach, V2P, draws inspiration from two distinct areas. To create attention "valleys" and suppress background noise, we adopt attention suppression techniques that penalize focus outside the target region (Li et al., 2018). To form a sharp "peak" at the target’s center, we are inspired by both Fitts’ Law from Human-Computer Interaction (HCI) (MacKenzie, 1992) and the common practice of using Gaussian heatmaps in localization tasks like pose estimation (Fitts, 1954). To our knowledge, our work is the first to synergistically combine background suppression with center-focused peak modeling to simulate the human pattern of interaction with the UI elements.
@@ 3 Method
We introduce Valley-to-Peak (V2P), a method that reshapes the model’s attention landscape to mimic human focus patterns for precise GUI grounding. It achieves this through two synergistic components:
  • Suppression Attention Valley Constraint: Penalizes attention on irrelevant regions to form low-attention "valleys," effectively suppressing background distractions.
  • Fitts-Gaussian Peak Modeling: Models interaction likelihood with a size-adaptive 2D Gaussian, creating a sharp attention "peak" at the target’s most actionable center.
By jointly optimizing these objectives, V2P produces a continuous, spatially-aware attention map that overcomes the limitations of rigid, uniform labels used in prior work. Below, we first outline the overall architecture (Sec. 3.1), then detail the Suppression Attention (Sec. 3.2) and Fitts-Gaussian Peak Modeling (Sec. 3.3) components.
@@ 3.1 Model Architecture Overview
We build upon GUI-Actor (Wu et al., 2025), a coordinate-free visual grounding framework that localizes GUI actions through attention rather than coordinate regression. Given a screenshot $I$ and an instruction $q$, the model introduces a special token <ACTOR> in the output sequence as a contextual anchor. The final-layer hidden state of <ACTOR>, denoted $h_{\mathtt{<ACTOR>}}$, is used to compute action attention over image patch features $\{v_{1},\dots,v_{M}\}$ extracted by the vision encoder.
@@
$\tilde{v}_{1},\dots,\tilde{v}_{M}=\text{SelfAttn}(v_{1},\dots,v_{M})$ (1)
yielding contextualized representations. These are projected into a shared embedding space with $h_{\mathtt{<ACTOR>}}$ via separate MLPs:
$z$ $=\text{MLP}_{T}(h_{\mathtt{<ACTOR>}}),$ (2)
@@
$\begin{split}\alpha_{i}&=\frac{z^{\top}z_{i}}{\sqrt{d}}\\ a_{i}&=\frac{\exp(\alpha_{i})}{\sum_{j=1}^{M}\exp(\alpha_{j})}\end{split}$ (4)
@@ 3.2 Suppression Attention Constraint for Distraction Mitigation
$\mathcal{G}=\left\{i\in\{1,\dots,M\}\mid R_{i}\cap b=\emptyset\right\}$ (5)
We define the attention loss as the total attention mass over these irrelevant regions:
$\mathcal{L}_{\text{Attn}}=\sum_{i\in\mathcal{G}}a_{i}$ (6)
To better understand the theoretical foundation of this constraint, we analyze the gradient dynamics of attention weights. For the target patch $k$ with attention weight $A_{k}=\text{softmax}(s_{k})$, the gradient with respect to any non-target patch logit $s_{i}$ is:
$\begin{split}w_{i}&=\frac{\partial A_{k}}{\partial s_{i}}=\frac{\partial\text{softmax}(s_{k})}{\partial s_{i}}\\ &=-\frac{e^{s_{k}}e^{s_{i}}}{(\sum_{i}^{M}e^{s_{i}})^{2}}=-A_{k}A_{i}<0\quad(i\neq k).\end{split}$ (7)
This gradient analysis reveals that any increase in attention logits $s_{i}$ for non-target patches negatively impacts the target attention $A_{k}$. The magnitude $|w_{i}|=A_{k}A_{i}$ quantifies this negative influence: larger values indicate that even small increases in attention to patch $i$ will cause rapid degradation in target attention $A_{k}$. This theoretical insight naturally motivates using $|w_{i}|$ as a weighting factor in our suppression loss, providing stronger penalties for patches that pose greater threats to target attention focus. And we have the suppression attention loss combined with gradient weight as:
$\mathcal{L}_{\text{Sup\_Attn}}=\sum_{i\in\mathcal{G}}w_{i}a_{i}$ (8)
@@ 3.3 Fitts-Gaussian Peak Modeling for Center-Focused Grounding
Our supervision strategy is inspired by Fitts’ Law (MacKenzie, 1992; Fitts, 1954), which reveals that click probability peaks at the center of an UI element and decays toward its edges, closely following a Gaussian distribution. We encode this behavior with Fitts-Gaussian Peak Modeling to guide the model’s focus in line with observed human interaction.
Specifically, we model the ideal attention distribution as a 2D Gaussian density centered at the centroid of the ground-truth bounding box $b=[x_{1},y_{1},x_{2},y_{2}]$:
$\mu=(c_{x},c_{y})=\left(\frac{x_{1}+x_{2}}{2},\frac{y_{1}+y_{2}}{2}\right)$ (9)
@@
$\sigma_{x}=\frac{w}{\sigma_{\text{factor}}},\quad\sigma_{y}=\frac{h}{\sigma_{\text{factor}}}$ (10)
@@
$\begin{split}y_{i}&=\left[\Phi(x^{i}_{\max};c_{x},\sigma_{x})-\Phi(x^{i}_{\min};c_{x},\sigma_{x})\right]\\ &\quad\cdot\left[\Phi(y^{i}_{\max};c_{y},\sigma_{y})-\Phi(y^{i}_{\min};c_{y},\sigma_{y})\right]\end{split}$ (12)
@@
To supervise the model’s predicted attention distribution $\{a_{i}\}$, we adopt the action attention loss from GUI-Actor (Wu et al., 2025), using the Kullback-Leibler (KL) divergence to measure the discrepancy between the target $p$ and prediction $a$:
$\begin{split}\mathcal{L}_{\text{Action\_Attn}}&=\sum_{i=1}^{M}p_{i}\log\frac{p_{i}}{a_{i}},\\ p_{i}&=\frac{y_{i}}{\sum_{j=1}^{M}y_{j}+\epsilon},\\ &\quad i=1,\ldots,M\end{split}$ (13)